# Channel Twinning: An Enabler for Next-Generation Ubiquitous Wireless Connectivity

### v1

@@ Abstract
The emerging concept of channel twinning (CT) has great potential to become a key enabler of ubiquitous connectivity in next-generation (xG) wireless systems. By fusing multimodal sensor data, CT advocates a high-fidelity and low-overhead channel acquisition paradigm, which is promising to provide accurate channel prediction in cross-domain and high-mobility scenarios of ubiquitous xG networks. However, the current literature lacks a universal CT architecture to address the challenges of heterogeneous scenarios, data, and resources in xG networks, which hinders the widespread deployment and applications of CT. This article discusses a new modularized CT architecture to bridge the barriers to scene recognition, cooperative sensing, and decentralized training. Based on the modularized design of CT, universal channel modeling, multimodal cooperative sensing, and lightweight twin modeling are described. Moreover, this article provides a concise definition, technical features, and case studies of CT, followed by potential applications of CT-empowered ubiquitous connectivity and some issues requiring future investigations.
@@ I Introduction
Ubiquitous connectivity is regarded as an overarching and fundamental function of next-generation (xG) networks (e.g., 6G networks), as it underpins seamless wireless coverage and network resilience to natural disasters [1]. To have ubiquitous connectivity, advanced coverage extension solutions such as space-air-ground-sea-integrated network (SAGSIN), cell-free network, and reconfigurable intelligent surface (RIS) have been proposed [2]. The provisioning of zero coverage-hole networks has to deal with the challenges related to diversified wireless channels across heterogeneous networks and multiple frequency bands, thus rendering network planning and optimization more convoluted than ever. This necessitates attention to one of the fundamental challenges in wireless communication systems, namely, channel acquisition, which determines the practical performance limits of wireless systems and networks.
I-A Paradigm Shifts in Channel Acquisition
Recently, channel twinning (CT) has emerged as a new regime of channel acquisition that reflects the complex interactions of channels and physical environments. CT has received increasing attention because it subverts traditional channel acquisition in the following research paradigms.
I-A1 A shift of perspective from mathematical to environmental modeling
Mathematical models (statistical and deterministic models) cannot balance the accuracy and complexity [3]. Instead, CT aims to reveal the channel mapping from environmental attributes, and achieve an accurate channel acquisition with low complexity.
I-A2 A shift from unimodal to multimodal sensing
Traditional channel acquisition primarily relies on radio frequency (RF) sensing-based pilot training methods, accompanied by high overhead. In contrast, multimodal fusion is a pillar of CT, and it is promising to reduce or even eliminate RF pilots [4].
I-A3 A shift from physical-world to digital-world measurement
In xG networks, cross-domain and high-mobility environments can cause poor channel estimation accuracy. While channel prediction can apply to dynamic scenarios, it is often not environment-aware and is limited by the correlation model. By emulating the interactions between the channel and the physical world, CT can create real-time channel replicas in the digital world.
@@ I-A3 A shift from physical-world to digital-world measurement
Concept
Role
Execution
phase
@@
Method
Accuracy
Generalizability
Complexity
Environment
synchronization
Research
perspective
Chanel modeling Derive proper models for effective channel representation Before network deployment
Statistical
modeling
Low $\times$× Low $\times$× Mathematical representation
Deterministic
modeling
High $\times$× High $\times$×
Hybrid
modeling
Moderate $\times$× Moderate $\times$×
@@
Obtain current
channel parameters
After network
deployment
Pilot-based or
training-based
Moderate $\times$× Moderate $\times$×
Mathematical
representation
@@
Infer future
channel states
After network
deployment
Time series
analysis
Moderate $\times$× Moderate $\times$×
Mathematical
representation
Channel twinning Create the digital replica of the channel evolution Full life cycle Synthetic CT High $\times$× Low $\times$× Environment semantics
Transfer CT Moderate Moderate
Multiverse CT Multi-level High
@@ I-B Current Status of CT Studies
In the context of CT, some initial studies have been reported in [5, 6, 7, 8]. The term of DT channel (DTC) [^1] To emphasize that “channel twinning” is a technology, similar to channel modeling or channel sounding, we have not used the terminology “DTC”. appeared earlier in [5], but the concept has not been specified. Later, the potential applications of predicting channels via digital twin (DT) were highlighted in [6], although the DTC concept was not proposed. The authors of [7] mentioned that DTC should enable channel prediction but did not provide clear definitions. More recently, DTC has been defined as the mapping between channels and associated communication operations [8]. Therein, the feasibility of real-time DTC has been demonstrated experimentally. As an initial instantiation of the CT concept, Boston Twin [9] was developed to introduce a city-scale twinning system for wireless networks.
The above literature has focused primarily on beam prediction applications or only discusses the knowledge graph construction from a local and centralized perspective. However, deploying CT in xG networks faces the challenges of cross-domain modeling, distributed sensing, and lightweight training. A centralized strategy makes it hard to cater to heterogeneous resources in xG networks.
Therefore, in this article, a systematic discussion on the modularized architecture of ubiquitous CT is conducted to facilitate the organic integration of CT and ubiquitous sensing and intelligence. For a better understanding of CT, a concise definition and several clarifications regarding the general CT misconceptions are refined. This article also highlights the technical evolutions, the potential use cases for CT-empowered ubiquitous connectivity technologies, and several promising directions to spur future research.
Fig. 1: Conceptual framework of CT in xG networks with ubiquitous connectivity. Advanced ubiquitous connectivity scenarios give birth to diversified wireless channels across heterogeneous networks. All of these channels can be acquired in the CT system without the need for high-overhead channel estimation.
@@ II-A Concepts and Realization Principle of CT
In this article, CT is defined as a digital mirror construction of physical channels for network design, optimization, and evaluation without real deployment. Essentially, such a digital mirror is a mapping model that indicates the correspondence from the environment to the associated channel state information (CSI). Specifically, a site-specific CT attempts to find a mapper to approximate the RF propagation law. As shown in Fig. 1, given the environmental conditions, CT can acquire the CSI at any position.
@@
Regarding the realization of CT, current mainstream methods can be primarily divided into three categories: Spatial interpolation modeling, environment semantic modeling, and deep learning (DL) modeling.
II-A1 Spatial Interpolation Modeling
Spatial interpolation [2] only utilizes channel measurements within the region of interest (RoI) to predict the channel of any location.
II-A2 Environment Semantic Modeling
Environment semantic-based methods, such as spatial loss field and virtual obstacle model [10], estimate the fading distribution of the entire RoI by observations.
II-A3 Deep Learning Modeling
@@
In DL methods, differentiable ray tracing (DRT) is gaining popularity as a promising DT enabler [11]. The DRT principle is to learn the electromagnetic (EM) properties of involved materials by modeling the interaction between ray paths and the environment.
@@
In practice, suitable methods should be selected according to applications and dataset sizes. From the perspective of implementation complexity, CT deployment should undergo three evolution stages: i) Synthetic CT [7], which is deployed in a quasi-static environment to generate synthetic data that approximates the measurements in the RoI; ii) Transfer CT [12], which constructs self-adaptive CT based on knowledge transfer to cope with unforeseen scenarios; and iii) Multiverse CT [13], which customizes different twins for applications with varying computing power and delay constraints. Each twin model offers multiple fidelity levels with different processing delays.
II-B Common Misconceptions and Three Clarifications
@@
To elaborate on the concept of CT comprehensively, we discuss the CT model illustrated in Fig. 1 from inputs to outputs. Several basic features are highlighted as follows: a) context-awareness: Instead of merely inputting received pilots or feedback by RF sensing methods, CT integrates contextual information from multimodal sensors (e.g., camera, GPS and Radar) to promptly adapt to environmental dynamics; b) high-fidelity: CT allows for high-fidelity emulation of complex interactions between the environment and the network. Thanks to three-dimensional (3D) map-aided methods, e.g., ray-tracing, spatial consistency [1] and multi-hop propagation can be simulated; c) self-synchronization: CT can model the whole life cycle of channel behaviors in the wireless environment for real-time emulation; d) expertise-driven: EM and channel modeling expertise are needed to design CT models. This helps produce compact DL structures and improve interpretability when specific channel characteristics, e.g., signal-to-noise ratio (SNR), are desired to be output; and e) region-specific: Since the CT model is trained on measurements from a specific site’s coverage region, the DTC must be region-dependent.
@@
Since CT is an emerging technology, there are several common misconceptions that need to be clarified.
Clarification 1: Do widely-used statistical channel modeling techniques belong to CT technology?
Following the CT concept, a DTC refers to a digital mirror of the channel response in a given environment, whereas standard statistical models like Rayleigh and Rician, cannot match specific environments. Moreover, 3GPP urban microcell (UMi) and urban macrocell (UMa) models only consider macroscopic environmental attributes (e.g., urban or rural) but ignore the specific environmental factors. In this sense, statistical models may not create the real DTCs.
@@
According to specific application requirements, one can flexibly choose different CT training methods, e.g., spatial interpolation, environment semantic, and DL methods.
@@
CT is either explicitly or implicitly context-aware. Only position-based methods construct the channels under the simple line-of-sight (LoS) assumption that ignores the environmental factors. However, blockages and scatters may cause non-LoS (NLoS) cases. This shows the significance of context awareness, and that position-based channel recovery cannot compete with CT.
@@ III A Modularized Design of CT Architecture
    Since xG networks are anticipated to involve massive cross-domain links across space/frequency/time-domains, how can we make the channel modeling of CT applicable to any scenario?
@@
In light of these questions, we present a modularized architecture to discuss the design considerations for CT deployment in xG networks. The overall CT architecture comprises three modules of universal channel modeling, multimodal cooperative sensing, and lightweight decentralized CT training, as shown in Fig. 2.
Fig. 2: The overall modularized CT architecture. The universal channel modeling module provides information about the sensing targets and prediction tasks of interest for the other two modules. The channel modeling module can generate synthetic data, which can be leveraged to pretrain the CT model in the lightweight twin module. Then, the CT model can be fine-tuned by feeding sensing measurements from the multimodal cooperative sensing module.
III-A Universal Channel Modeling Module
Various scenarios in xG networks give rise to various types of channels. The characteristics of these channels differ significantly from each other. Consequently, a major challenge that arises for CT deployment is how to integrate the diverse channel characteristics across scenarios into a unified model. To this end, a universal channel modeling module tailored for ubiquitous CT is needed.
Generally, a channel consists of large-scale fading (LSF) and small-scale fading (SSF) components, each with different characteristics in various scenarios. Therefore, the channel components associated with the scenario and working band should be identified at first. For instance, high-mobility scenarios, such as Internet of Vehicles (IoVs) and unmanned aerial vehicle (UAV) networks, should consider non-stationary channels. Gas absorption and blocking effects should be considered in millimeter-wave (mmWave) and Terahertz (THz) networks, while visible light communication (VLC) channels only involve LSF. Extremely large-scale, multiple-input-multiple-output (MIMO) channels should consider spherical wavefront and spatial non-stationarity [1].
To effectively deploy CT, we can quickly determine the channel components by following two steps: 1) LSF channel construction by a static 3D map; and 2) SSF channel prediction by capturing dynamic activities.
@@ III-B Multimodal Cooperative Sensing Module
The traditional channel measurement sources come primarily from RF reception and sensing. Such measurement is limited by the sensing range and accuracy of the single-site ISAC scheme. For RF sensing, a multi-site cooperative sensing scheme can be preferable, which can enhance signal acquisition by virtue of ubiquitous connectivity. In this scheme, non-terrestrial and terrestrial base stations (BSs) can sense any position through ISAC technology. In addition, users can also track targets by the dedicated positioning reference signals defined in 5G-NR R17 [14].
However, cooperative sensing requires precise space-time registration, which is crucial for positioning data acquisition and continuous updates of CT systems. Non-RF sensing involves more heterogeneous multimodal data than RF sensing. In this sense, the challenge of M2SDF lies in how to deal with the data heterogeneity at the CT server. To extract propagation environment semantics, semantic alignment and completion [8] techniques may be invoked.
Fig. 3: Training architecture of CT: (a) Centralized learning-based CT model construction, where dedicated measurement anchors collect data and upload data to the CT server for training; and (b) FL-based CT model construction, where registered device clients use their local data to train the model and only upload trained parameters for model aggregation at the CT server.

### v3

# Advancing Ubiquitous Wireless Connectivity through Channel Twinning

@@ Abstract
As an emerging trend in channel acquisition (CA), the concept of channel twinning (CT) has been proposed as a powerful enabler of ubiquitous connectivity in next-generation (xG) wireless systems. By fusing multimodal sensor data, CT advocates a high-fidelity and low-overhead CA paradigm, which is promising to provide accurate channel prediction in cross-domain and high-mobility scenarios of ubiquitous xG networks. However, existing literature lacks a universal CT architecture to address the challenges of heterogeneous scenarios, data, and resources in xG networks, which hinders the widespread deployment and applications of CT. This article discusses a new modularized CT architecture to bridge scene recognition, cooperative sensing, and decentralized training, comprising versatile model configuration, multimodal cooperative sensing, and lightweight twin modeling modules. Additionally, this article presents a detailed concept, technical features, and case studies of CT, outlines mainstream trends of realization methods, followed by potential applications of CT-empowered ubiquitous connectivity, and issues requiring future investigations.
@@ I Introduction
Ubiquitous connectivity is regarded as an overarching and fundamental function of next-generation (xG) networks (e.g., 6G networks), as it underpins seamless wireless coverage and network resilience to natural disasters [1]. To have ubiquitous connectivity, advanced coverage extension solutions such as space-air-ground-sea-integrated network (SAGSIN), cell-free network, and reconfigurable intelligent surface (RIS) have been proposed [2]. The provisioning of zero coverage-hole networks has to deal with challenges related to diversified wireless channels across heterogeneous networks and multiple frequency bands, thus rendering network planning and optimization more convoluted than ever. This necessitates attention to one of the fundamental challenges in wireless communication systems: channel acquisition (CA), which determines the performance limits of wireless networks.
I-A Advances in Channel Acquisition
Recently, channel twinning (CT) has emerged as an innovative approach to CA that captures the intricate interactions between channels and physical environments. CT has received increasing attention for its ability to enhance traditional CA in the following ways:
I-A1 From mathematical to environmental modeling
It can be difficult for mathematical models (statistical and deterministic models) to strike a balance between channel prediction accuracy and computational complexity [3]. Instead, CT aims to reveal the channel mapping from environmental attributes and achieve an accurate CA with low complexity.
I-A2 From unimodal to multimodal sensing
Traditional CA primarily relies on radio frequency (RF) sensing-based pilot training methods, accompanied by high overhead. In contrast, multimodal fusion [4, 5] is a pillar of CT, and it is promising to reduce or even eliminate RF pilots.
I-A3 From physical-world to enhanced digital-world processing
Traditional channel modeling or channel estimation methods often rely on channel assumptions to fit the channel model solely based on measurements. CT exploits measurements to extract the environment semantics or interpolate the channel measurements at any location, leading to a channel knowledge model that better reflects real-world environments. On the other hand, CT has the potential to significantly reduce the dependency on expensive real-world measurements once the CT model is constructed.
@@ I-A3 From physical-world to enhanced digital-world processing
{tblr}
width = colspec = Q[120]Q[140]Q[160]Q[160]Q[120]Q[180], cell15 = c=20.3, cell21 = r=3, cell25 = c=2,r=30.3, cell43 = c=20.3, cell55 = c=20.3, cell65 = c=20.3, cell71 = r=3, cell75 = r=3, vlines, hline1-2,5-7,10 = -, hline3-4 = 2-4, hline8-9 = 2-4,6, Concept & Modeling Method Pros Cons Role and Application
Chanel
@@
Statistical Versatility Limited accuracy • Role: Derive effective model representation
• Application: Pre-deployment
Deterministic High accuracy Time-consuming
Hybrid Balance accuracy and complexity
@@
Pilot-based Stable performance • High pilot load
• Outdated CSI • Role: Obtain current CSI
• Application: Post-deployment
@@
Time series analysis Proactive management Error accumulation • Role: Infer future channel states
• Application: Post-deployment
Channel
twinning
Spatial interpolation • Simple implementation
• Low data dependency • Limited accuracy
• Poor scalability • Role: Create the digital replica of the channel evolution
• Application: Full life cycle Stable scenarios with smooth channel changes
Environment semantics Enhanced modeling accuracy Expensive environment data collection Scenarios with significant environment structure effects
Deep learning • Automatic learning
• No explicit assumptions • Need large data
• Hyper-parameter tuning Highly dynamic and complex scenarios
@@ I-B Current Status of CT Studies
In the context of CT, some initial studies have been reported in [6, 7, 8]. The term of digital twin channel (DTC) [^1] To emphasize that “channel twinning” is a technology, similar to channel modeling or channel sounding, we have not used the terminology “DTC”. appeared earlier in [6], but the concept has not been specified. Later, the potential of channel recreation on digital twin (DT) was revealed in [7], although the DTC concept was yet to be proposed. More recently, DTC has been defined as the mapping between channels and associated communication operations [8]. Therein, the radio environment knowledge pool (REKP) was proposed to support DTC.
The above literature has focused primarily on channel knowledge map (CKM) applications [2] or only discusses the REKP construction from a local and centralized perspective. While providing a framework for reflecting propagation environment via channel fingerprints, CKM emphasizes static mappings. REKP [8] is a knowledge semantic database that stores and manages the feature relationships of radio environments and channels. However, it is not tailored to specific scenarios or environments. In xG networks, deploying CT faces the challenges of cross-domain modeling, distributed sensing, and lightweight training. A centralized strategy makes it hard to cater to heterogeneous resources.
In this article, a systematic discussion on the modularized architecture of ubiquitous CT is conducted to facilitate the organic integration of CT and ubiquitous sensing and intelligence. For a better understanding of CT, a concise definition and several clarifications regarding the general CT misconceptions are refined. This article also highlights the technical evolutions, the potential use cases for CT-empowered ubiquitous connectivity technologies, and several promising directions to spur future research.
@@ II-A Concepts and Realization Principle of CT
CT is defined as a digital mirror construction of physical channels for network design, optimization, and evaluation without real deployment. In essence, such a digital mirror is a mapping model that indicates the correspondence from the environment to various aspects of associated channel behaviors, such as channel state information (CSI), channel path angle, or channel gain, depending on the specific task. Specifically, a site-specific CT aims to simulate the RF propagation law. As shown in Fig. 1, CT can acquire CSI at any position based on the environmental conditions.
Fig. 1: Conceptual framework of CT in xG networks and its modularized CT architecture. The versatile model configuration module provides information for measurements and model design to the other two modules. It generates synthetic data to pretrain the CT model in the lightweight twin module. Then, the CT model can be fine-tuned by feeding data from the multimodal cooperative sensing module.
@@
Current methods for realizing CT fall into three main categories: Spatial interpolation (SI) modeling, environment semantics (ES) modeling, and deep learning (DL) modeling.
II-A1 SI
SI leverages the spatial correlations of sampling channel measurements within the region of interest (RoI) to predict the channel of unobserved locations. Classic SI methods include inverse distance weighting, Kriging, etc., commonly used for constructing CKM [2].
II-A2 ES
Utilizing ES features (e.g., building layout, terrain type, and obstacles), ES modeling can help accurately build channel models, such as the virtual obstacle model [9] and REKP [8].
II-A3 DL
@@
DL methods apply advanced deep networks to learn the complex mapping between various environmental factors affecting the channels. In DL methods, differentiable ray tracing (DRT) is gaining popularity as a promising DT enabler [10]. The DRT principle is to learn the electromagnetic (EM) properties of involved materials by modeling the interaction between ray paths and the environment.
@@
In practice, suitable methods should be selected according to applications and dataset sizes. From the perspective of implementation complexity, CT deployment should undergo three evolution stages: i) Synthetic CT [7], which is deployed in a quasi-static environment to generate synthetic data that approximates the measurements in the RoI; ii) Transfer CT [11], which constructs self-adaptive CT based on knowledge transfer to cope with new scenarios; and iii) Multiverse CT [12], which customizes different twins to offer multiple fidelity levels for applications with diverse computing power and delay constraints.
II-B Common Misconceptions and Four Clarifications
@@
To elaborate on the concept of CT, we discuss the CT model illustrated in Fig. 1 from inputs to outputs. Several basic features are highlighted as follows: a) context-awareness: Instead of merely using received pilots or feedback, CT integrates contextual information from multimodal sensors (e.g., camera, GPS, and Radar) to learn environmental dynamics; b) high-fidelity: Thanks to three-dimensional (3D) map-aided methods [9, 10], e.g., ray-tracing, CT allows for high-fidelity emulation of complex interactions between the environment and the network; c) self-synchronization: CT can model the whole life cycle of channel behaviors in the environment for real-time emulation; d) expertise-driven: EM and channel modeling expertise help design compact DL structures and improve interpretability when specific channel properties, e.g., signal-to-noise ratio (SNR), are desired to be output; and e) region-specific: Since the CT model is trained on measurements from a specific site’s coverage region, the DTC must be region-dependent.
@@
Since CT is an emerging technology, several common misconceptions need to be clarified.
Clarification 1: Do statistical and deterministic modeling belong to CT technology?
CT is designed to create a digital mirror of the channel response in a given environment, whereas standard statistical models like Rayleigh and Rician cannot match specific environments. Besides, 3GPP urban microcell (UMi) and urban macrocell (UMa) models only consider macroscopic environmental attributes (e.g., urban or rural) but ignore the specific environmental factors [3]. In this sense, statistical models may not reflect the CT features. Deterministic models, such as ray tracing, is a physics-based model that predicts channel by simulating ray propagation in specific environments and scene geometries, as well as complex EM interactions with surrounding objects. On the other hand, CT is a mapping model-based CA method that seeks to efficiently mimic channel behaviors with no need for exhaustive recursive tracing. As a preliminary attempt at CT, Sionna is ten times faster than traditional ray-tracing like Wireless InSite (WI) [10]. Conventional channel modeling methods lay a solid foundation for understanding the physical behaviors of waves, while CT can adaptively respond to environmental changes through its self-synchronization capability, as discussed in Section II-B. Combining ray tracing with CT can further contribute to advanced channel modeling, and potentially bring real-time ray tracing to reality.
@@
According to specific application requirements, one can flexibly choose different CT training methods, e.g., SI, ES, and DL modeling methods.
@@
CT is either explicitly or implicitly context-aware. Only position-based methods construct the channels under the simple line-of-sight (LoS) assumption. However, blockages and scatters may cause non-LoS (NLoS) cases. This shows the significance of context awareness, and that position-based channel recovery cannot compete with CT.
Clarification 4: What are the conceptual differences between DT, CT, and DTC?
A DT is a virtual replica of a physical entity (e.g., a network or device). DTC develops an REKP to help understand the interaction between channels and environmental characteristics across extensive scenarios [8]. The REKP framework updates the path loss feature relationship graphs for different receiver locations through real-time acquisition of environmental features (e.g., position, blockage, and distance). However, these updates rely on static environmental factors within localized time intervals, rather than comprehensive dynamic modeling of channel evolution over time. In contrast, CT focuses on learning both static attributes and the temporal evolution of channels, emphasizing the twinning process in a specific region.
@@ III A Modularized Design of CT Architecture
    Since xG networks are anticipated to involve massive cross-domain links across space/frequency/time domains, how can we streamline the CT model deployment and configuration process for diverse scenarios and channel types?
@@
In light of these questions, we present a modularized architecture to discuss the design considerations for CT deployment in xG networks. The overall CT architecture comprises three modules of versatile model configuration, multimodal cooperative sensing, and lightweight decentralized CT training, as shown in Fig. 1.
III-A Versatile Model Configuration Module
Various scenarios in xG networks give rise to various types of channels. The characteristics of these channels differ significantly from each other. Consequently, a major challenge that arises for CT deployment is how to integrate the diverse channel characteristics across scenarios to facilitate the automated construction of suitable CT models. To this end, a versatile model configuration module tailored for ubiquitous CT is needed.
Generally, a channel consists of large-scale fading (LSF) and small-scale fading (SSF) components, each with different characteristics in various scenarios. Therefore, the channel components associated with the scenario and working band should be identified first. For instance, high-mobility scenarios, such as Internet of Vehicles (IoV) and unmanned aerial vehicle (UAV) networks, should consider non-stationary channels. Gas absorption and blocking effects should be considered in millimeter-wave (mmWave) and Terahertz (THz) networks, while visible light communication (VLC) channels only involve LSF. Extremely large-scale, multiple-input-multiple-output (MIMO) channels should consider spherical wavefront and spatial non-stationarity [1].
To effectively deploy CT, we can quickly determine the channel components by following two steps: 1) LSF channel construction by a static 3D map; and 2) SSF channel prediction by capturing dynamic activities. The versatile model configuration module allows developers to easily configure the CT model by selecting from various model settings and adjusting parameters to suit specific task needs. Additionally, this module provides supplementary synthetic data to assist with CT model training, rather than engaging in traditional channel modeling.
@@ III-B Multimodal Cooperative Sensing Module
The traditional channel measurements come primarily from RF reception and sensing. Such measurement is limited by the sensing range and accuracy of the single-site ISAC scheme. For RF sensing, a multi-site cooperative sensing scheme is preferable, which can enhance signal acquisition by ubiquitous connectivity. In this scheme, non-terrestrial and terrestrial base stations (BSs) can sense any position through ISAC. In addition, users can also track targets by the dedicated positioning reference signals defined in 5G-NR R17 [13].
However, cooperative sensing requires precise space-time registration, which is crucial for positioning data acquisition and continuous updates of CT systems. Non-RF sensing involves more heterogeneous multimodal data than RF sensing. In this sense, the challenge of M2SDF lies in how to deal with the data heterogeneity at the CT server. ES plays a crucial role in enabling CT by providing detailed information about the physical environment. The ES implementation in the CT framework consists of the following steps: 1) Extracting physical structure features of the environment that impact signal paths, such as walls, trees, and other obstructions [9]; 2) classifying and encoding these environment features to construct channel semantic information. To extract propagation ESs, semantic alignment, and completion [8] techniques may be invoked; and 3) utilizing semantic features to model the interactions between network actions and specific tasks. Thus, optimization algorithms or deep networks can be designed to improve the accuracy of CA, adapt to dynamic scenarios, or enhance communication performance.
Fig. 2: Training architecture of CT: (a) Centralized learning-based CT, where dedicated measurement anchors collect data and upload it to the CT server for training; and (b) FL-based CT, where registered device clients use their local data to train the model and only upload trained parameters for model aggregation at the CT server.